A park has two sprinkers that are used to fill a fountain.One sprinker can fill the fountain in 3 hours and the second sprinker can fill the fountain in 6 hours. How long will it take to fill the fountain with both sprinkers operating.
Found 2 solutions by chitra, tanimachatterjee:
Answer by chitra(359)

It is given that the sprinkler can fill the fountain in 3 hours.

So in 1 hr it fills 1/3 of the fountain.

It is also given that another sprinkler can fill the fountain in 6 hours.

So in 1 hr it fills 1/6 of the fountain.

So, the total time taken to fill the fountain by both the spriklers is:

1/3 + 1/6 = 3/6

this is equal to 1/2

So, the total time taken to fill the fountain by both the sprinklers is 2 hours.


Happy Calculating...

Regards
Chitra

Answer by tanimachatterjee(60)

One sprinker can fill the fountain in 3 hours
therefore in 1 hr. it can fill 1/3 portion of the fountain
second sprinker can fill the fountain in 6 hrs.
in1 hr it can fill 1/6 portion
when both sprinkers operating
in 1 hr. they can 1/3 + 1/6 = 1/2 portion of the fountain
so they will take 2 hrs to fill the tab